Circular No. 9131

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 F3 (SCOTTI) J. V. Scotti, Lunar and Planetary Laboratory, reports his discovery of

another comet, described as slightly diffuse with a tail on 0.9-m Spacewatch reflector CCD mosaic images taken on Mar. 20 (discovery observation tabulated below); co-added images taken on Mar. 20.3 UT with the Spacewatch 1.8-m reflector (6.8 min exposure time) show a coma diameter of 5′′ with a narrow tail that extends to at least 1′.06 (with a possible faint extension out to 1′.47) in p.a. 289^{o}. Scotti adds that nine co-added images with the latter telescope on Mar. 21.3 show a 7′′ coma with a long, narrow tail extending at least to 1′.9 in p.a. 289^{o}, and possibly as far as 3′.0 in p.a. 288^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have also commented on the object’s cometary appearance. W. H. Ryan (Magdalena Ridge 2.4-m reflector, Mar. 22.3) finds that his R-band images show a coma and a thin tail in p.a. ∼ 290^{o}. M. Busch and D. Koschny measaured a tail ∼ 20′′ long in p.a. 280^{o} on images taken on Mar. 22.96 by D. Abreu with the 1.0-m f/4.4 reflector of the ESA Optical Ground Station at Tenerife. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m reflector, Mar. 23.0) finds a coma diameter of ∼ 8′′ and a straight, thin tail 30′′ long in p.a. 287^{o}.

2010 UT α_{2000} δ_{2000} Mag. Observer
Mar. 20.18761 11^{h}36^{m}24.^{s}37 +4^{o}19′05′′.2 19.8 Scotti

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-F67.

T = 2010 Nov. 2.199 TT  ω = 42.^{o}477
Ω = 157.151 }2000.0
q = 5.23577 AU  i = 4.572

COMET C/2009 O2 (CATALINA) H. Sato (Ota-ku, Tokyo, Japan, 0.30-m reflector) finds a very diffuse

coma (3′ × 5′) with no condensation on images taken on Mar. 12.5 and 13.5 UT, suggesting an apparent disintegration of this comet. L. Buzzi, Varese, Italy, reports that images with taken on Mar. 13.18 (0.60-m reflector) show no clear central condensation (contrary to his images from Mar. 5.18). H. Mikuž, Crni Vrh Observatory, finds that R-band images taken on Mar. 8.16 with a 0.60-m telescope show a coma of mag 16 without distinct central condensation and a fan-like tail ∼ 1′.5 long in p.a. ∼ 300^{o}, while images from Mar. 14.17 show a fading to mag ∼ 17.

2010 March 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green